Slotkin Products purchased a machine for $39,000 on July 1, 2014. The company intends to depreciate it over 8 years using the double-declining balance method. Salvage value is $3,000. Depreciation for 2015 to the closest dollar is
Answer #1

Straight line rate = (100/8)% = 12.5%

Double declining rate = (2*12.5)% = 25%

Depreciation for year 2014 = 39,000 * 25% * 6/12 = 4875

Depreciation for 2015 = (39,000-4875) * 25% = 8531

Comment if you face any issues